Andy is twice as old as Kate. In 6 years, their ages will total 60. How old is Andy now?
Answer by stanbon(75887)

Andy is twice as old as Kate. In 6 years, their ages will total 60.
How old is Andy now?
-----------
A = 2K
A+6 + K+6 = 60
--------
Substitute for "A" and solvee for "K":
2K+6 + K+6 = 60
3K + 12 = 60
3k = 48
K = 16 yrs (Kate's age now)
---
A = 2K = 2*16 = 32 yrs (Andy's age now)
------------
Cheers
Stan H.
-----------